OpenStudy (anonymous):

2 divided by x+2 plus 1 divided by 5 is equal to 6 divided by x +5....... find x :)

OpenStudy (anonymous):

a.) x=0 b.)x= -13 c.) x=13 d.) x=0, x=13

OpenStudy (anonymous):

-50/29

OpenStudy (gw2011):

The answer is (d) x=0, x=13

OpenStudy (anonymous):

oops didnt see the +5

OpenStudy (gw2011):

Here is how I calculated the answer to be (d): 2/(x+2) + 1/5 = 6/(x+5) Multiply each term by (x+2) to eliminate the denominator 2 + (x+2)/5 = (6)(x+2)/(x+5) Multiply each term by (x+5) for the same reason 2(x+5) + (x+2)(x+5)/5 = 6(x+2) Multiply each term by 5 10(x+5) + (x+2)(x+5) = 30(x+2) 10x + 50 + x^2 + 7x + 10 = 30x + 60 x^2 +17x +60 = 30x + 60 x^2 + 17x - 30x = 60 - 60 x^2 - 13x = 0 x(x-13) = 0 x = 0 This is one of the solutions x-13 = 0 x = 13 This is the other solution Therefore, x=0 and x=13 which is the (d) choice